A swimming pool is 20yds. by 10 yds and is 5 feet deep. How many square feet are in the surface area of the sides and the bottom of the pool.
Answer by rfer(16322)

you have 5 areas to calculate
two sides are equal and two lengths are equal
20 yds=60 ft
10 yrs=30 ft
2(5*30)+2(5*60)+(30*60)=
300+600+1800=2700 sq ft total